OpenStudy (lgbasallote):

Find the area bounded by the curves \(y = e^x\) and \(y = x^2 - 2\) from x = -1 to x = 1

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}e^x - (x^2-4)dx = (e^x -\frac{x^3}{3} +4x)|_{-1}^{1}\] =\[(e - \frac{1}{3}+4)-(\frac{1}{e}+\frac{1}{3}-4) = 8-\frac{1}{e}+e\]

OpenStudy (lgbasallote):

\[\large e - \frac{1}{3} + 4 - \frac{1}{e} - \frac{1}{3} + 4\] then it should be \[\large e - \frac{1}{2} - \frac{2}{3} + 8\] @SmoothMath

OpenStudy (lgbasallote):

although i think it should be 2 not 4

OpenStudy (lgbasallote):

i dont recall the question well lol

OpenStudy (lgbasallote):

uhh that was supposed to be \[\large e - \frac{1}{e} - \frac{2}{3} + 8\]

OpenStudy (anonymous):

Ah, yes. Correct. And why would it be 2? 4x is the anti-derivative of 4.

OpenStudy (lgbasallote):

no...i meant the question should be the are bounded by \(y = e^x\) and \(y =x^2 -2\) not 4...anyway...ill just see the process here...what comes next?

OpenStudy (anonymous):

I mean... nothing really. You can rewrite that if you like though.

OpenStudy (lgbasallote):

so here was my solution.... \[\int_{-1}^1 e^x -(x^2 - 2)dx\] \[\int_{-1}^1 e^x - x^2 + 2\] \[[e^x|_{-1}^1 - [\frac{x^3}{3}|_{-1}^1 + [2x|_{-1}^1\] \[[e - \frac{1}{e}] - [\frac{1}{3} - (-\frac{1}{3})] + [2 - (-2)]\] \[(\frac{e^2 - 1}{e}) - \frac{2}{3} + 4\] \[\frac{3(e^2 -1) - 2e + 12e}{3e} = \frac{3(e^2 - 1) + 10e}{3e}\] that right?? @smoothmath ?

OpenStudy (anonymous):

Quite right. =) Yeah, I changed it to 4 for some reason. Silly me.

OpenStudy (lgbasallote):

no it was originally 4 lol =))

OpenStudy (lgbasallote):

it seems that was the question

OpenStudy (anonymous):

or you could simplify that down to \[\frac{3e^2 + 10e - 3}{3e}\]

OpenStudy (anonymous):

Oh, hahaha

OpenStudy (lgbasallote):

thanks...that was a great reassurance that i can actually pass the subject :DDDD